OpenStudy (anonymous):

Fact5or the polynomial 21x3-12x2

OpenStudy (anonymous):

Hint: Prime factorization of the terms gives \(3 \cdot 7\cdot x \cdot x\cdot x - 2\cdot 2 \cdot 3 \cdot x \cdot x\). What are the common factor(s) you can pull out (GCF)?